If ALICE is twice as old as BRENDA and BRENDA is half as old as CATHERINE, How old is each of them when their total ages are 50 ?
Found 2 solutions by Maths68, richwmiller:
Answer by Maths68(1474)

Let
Present age of Catherine = x
Present age of Brenda = x/2
Present age of Alice = 2*(x/2) = x
x+x/2+x=50
2x+x/2=50*2
Multiply by 2 both sides of the equation
4x+x=100
5x=100
5x/5=20
x=20
Present age of Catherine = x = 20 years old
Present age of Brenda = x/2 = 20/2 = 10 years old
Present age of Alice = 2*(x/2) = x = 20 years old

Answer by richwmiller(17219)

no similar problems
a=2b
b=c/2
c=2b
a+b+c=50
5b=50
b=10
a=20
c=20